# Towards a turnkey approach to unbiased Monte Carlo estimation of smooth functions of expectations

### v2

@@ 1.1. General setting
These conditions ensure that $\hat{f}$ is an unbiased estimator, see the following lemma (and Appendix A for a proof).
@@ 1.2. Motivation and contributions
We are particularly interested in the functions $f(x)=\log x$ and $f(x)=1/x$. The former arises in applications such as maximum likelihood estimation where unbiased estimates of the likelihood are available, for example using importance sampling. Using the sum estimator framework, these can then be converted into unbiased estimates of the log-likelihood (and its gradient). The function $f(x)=1/x$ arises in Bayesian and frequentist estimation of models with un-normalised likelihoods, that is, when the likelihood involves an intractable normalising constant.
Various unbiased estimators have already been proposed in the literature for such functions; see Section 1.3 for a review. However, in our experience, their performance tends to be very sensitive to ‘tuning’ parameters such as the expansion point $x_{0}$ and the distribution of $R$. This is problematic for users, in particular in situations where one needs to routinely compute unbiased estimates of $f(m)$ at many different $m$ values. For instance, in stochastic gradient methods, where at each iteration one needs to estimate the gradient of a different term; see Section 2.6 for details.
Thus, our motivation is to propose a general ‘turnkey’ approach to unbiased estimation which is both robust and easy to automate. Specifically, our contributions are as follows. We derive (Section 2) some results on the general properties of the randomly truncated Taylor-series based estimator, and use them to inform the choice of the truncation random variable $R$. We also discuss in detail the properties of the standard unbiased estimate (4) of $(m/x_{0}-1)^{k}$ and propose an alternative with potentially much smaller variance. In particular, our new estimate uses the whole collection of samples $X_{1},\ldots,X_{r}$ to construct the unbiased estimate $U_{r,k}$ of $(m/x_{0}-1)^{k}$ for $0<k\leq r$. We show the overall variance of the sum estimator will then decrease at a rate $\mathcal{O}\left(\log\mathbb{E}[R]/\mathbb{E}[R]\right)$. In contrast, the variance of the sum estimate with the standard choice (4) for the coefficient estimate does not diminish to zero as $\mathbb{E}[R]$ increases. This difference can be significant for some applications.
The second part of our study (Section 3) focuses on the specific functions $f(x)=\log x$ and $f(x)=1/x$. For these functions, we discuss how to choose tuning parameters ($x_{0}$ and the distribution of $R$) to ensure that the resulting estimator has finite – and hopefully small – variance.
@@ 1.4. Notations and standing assumptions
The expansion point $x_{0}\in\mathbb{R}\setminus\{0\}$ in (1) is fixed initially (and thus so are the $\gamma_{k}$’s), but we discuss how to choose $x_{0}$, for a particular $m$, later on in Section 3.
@@ 2.1. Variance of the general sum estimator
(5) $\mathbb{E}\left[\hat{f}\middle|R=r\right]=\sum_{k=0}^{\infty}\gamma_{k}\frac{\mathds{1}\{r\geq k\}}{\mathbb{P}(R\geq k)}\mathbb{E}(U_{r,k}|R=r)=\sum_{k=0}^{\infty}\gamma_{k}\frac{\mathds{1}\{r\geq k\}}{\mathbb{P}(R\geq k)}\left(\frac{m}{x_{0}}-1\right)^{k}.$
(See Appendix A.2 for the verification.)
One may decompose its variance (assuming it is finite) through the law of total variance,
@@
(See Appendix A.3 for the verification.)
The term $\operatorname{var}\left[\mathbb{E}[\hat{f}|R]\right]$ has a simple interpretation, namely, it is the variance of $\hat{f}$ in the ideal case where the $X_{i}$’s (and the $U_{r,k}$’s) have zero variance, that is, $U_{r,k}=(m/x_{0}-1)^{k}$. Hence, this variance component is only due to the random truncation of the Taylor expansion.
@@
For the function $f:\mathbb{R}\rightarrow\mathbb{R}$, argument $m$ and expansion point $x_{0}$, the series expansion in (1) holds. Furthermore, (i) for some $c>0$, $|\gamma_{k}|\leq c$ for all $k=0,1,\ldots$; (ii) $g^{k}=o(|\gamma_{k}|)$ for all $g\in(0,1)$; and (iii) $\beta_{0}:=|m/x_{0}-1|<1$.
@@
(See Appendix B.1 for a proof.) Assumption 1 is needed to ensure integrability (and thus $\mathbb{E}[\hat{f}|R]$ is well defined.) The condition $\beta_{0}<1$ (of Assumption 3) and corresponding restriction on $p$ ensures the variance is finite.
@@
Both our functions of interest fulfil Assumption 3: $\gamma_{k}=(-1)^{k-1}/k$ for $f(x)=\log x$, and $\gamma_{k}=(-1)^{k-1}$ for $f(x)=1/x$.
A function that does not fulfil Assumption 3 is $f(x)=\exp(x)$ as $\gamma_{k}$ decreases at a super-geometric rate. This makes it possible to consider for $R$ some distribution with tails that are lighter than geometric, for example, the Poisson distribution, which is why we do not considered it further, but again see Papaspiliopoulos, (2011) for a thorough treatment of this function.
Before moving on, we state the expression for the second term of decomposition (6), which we will study in the following sections.
@@ 2.2. Simple estimator
Note that, given $m$ and $\sigma^{2}$, it is always possible to find $x_{0}$ such that (10) holds; in fact the minimiser of $\beta$ with respect to $x_{0}$ verifies this condition.
What is unsatisfactory about this result is that the right hand side does not vanish as $p\rightarrow 0$. This is not because the bound is not tight, as shown by the following example (see Appendix B.3 for the derivation the limit of $\mathbb{E}\left[\operatorname{var}[\hat{f}^{\mathrm{S}}|R]\right]$ as $p\rightarrow 0$). Our numerical experiments suggest that this is true for other functions of interest.
@@ 2.3. Cycling estimator
$U_{r,k}^{\mathrm{C}}:=\frac{1}{r}\left[\prod_{i=1}^{k}\left(\frac{X_{i}}{x_{0}}-1\right)+\prod_{i=2}^{k+1}\left(\frac{X_{i}}{x_{0}}-1\right)+\ldots+\left(\frac{X_{r}}{x_{0}}-1\right)\prod_{i=1}^{k-1}\left(\frac{X_{i}}{x_{0}}-1\right)\right].$
@@
In other words, by taking $\mathbb{E}[R]\rightarrow+\infty$, the cycling estimate recovers, up to a log factor, the standard Monte Carlo rate, i.e. the variance is $\mathcal{O}\left(\log\{\mathbb{E}[R]\}/\mathbb{E}[R]\right)$. Thus, both terms of the variance in (6) now decrease as the mean of $R$ increases.
A direct corollary (through Chebyshev’s inequality) of the above result is that the cycling estimator converges in probability.
@@ 2.4. Minimum variance unbiased estimator
which has minimimum variance (among unbiased estimators of its expectation), see Hoeffding, (1948). In spite of the combinatorial number of terms, it is still possible to compute the estimator in $\mathcal{O}(r^{2})$ operations, using a recursion (see Appendix D). [^1] Pointed out to us by Sam Power in a personal communication. However, repeating our variance analysis for the MVUE estimator appears challenging, indeed due to the combinatorial number of terms in (12), and is left for further study. For the moment, we present a numerical comparison in Appendix D, which happens to show cycling and MVUE performing similarly.
@@ 2.5. Relative merits of the simple and cycling estimators
The cycling estimator makes it possible to estimate $f(m)$ as accurately as required, and enjoys (up to a log factor) the same error rate as Monte Carlo estimates. On the other hand, it is more expensive to compute than the simple estimator. Although both estimators require generating $R$ random variables, computing the $U_{R,k}$’s ($0\leq k\leq R$) for the cycling estimator requires $\mathcal{O}(R^{2})$ operations.
@@ 2.6. Unbiased estimators of gradients
and the corresponding ‘sum’ estimator (that is similar to (2) in construction) is
@@
where $R$ is a random variable taking values in $\{1,\ldots\}$; $W_{R,k}=\sum_{r=k}^{\infty}W_{r,k}\mathds{1}\{R=r\}$; $W_{r,k}$ (for $r\geq k$) is an unbiased estimate of $(m(\theta)/x_{0}-1)^{k-1}\nabla_{\theta}m(\theta)$; and each and every $W_{r,k}$ is independent of $R$.
As was stated for $\hat{f}$ in Lemma 2, the gradient estimate can be shown to be (integrable and) unbiased when the estimates $W_{r,k}$ satisfy similar conditions of Assumption 1, that is, with univariates $U_{r,k}$ therein replaced with $W_{r,k}$, and $\gamma_{k}$ replaced with $k\gamma_{k}$. (See Appendix A.1 for the proof of unbiasedness of $\hat{f}$ which can be easily adapted to the present gradient case.)
We can define a simple estimator for (13), denoted by $\widehat{\nabla_{\theta}f}^{\mathrm{S}}$, as follows:
@@
$W_{r,k}^{\mathrm{C}}:=\frac{1}{r}\Bigg{[}G_{k}\prod_{i=1}^{k-1}\left(\frac{X_{i}}{x_{0}}-1\right)+G_{k+1}\prod_{i=2}^{k}\left(\frac{X_{i}}{x_{0}}-1\right)\\ +\ldots+G_{k-1}\left(\frac{X_{r}}{x_{0}}-1\right)\prod_{i=1}^{k-2}\left(\frac{X_{i}}{x_{0}}-1\right)\Bigg{]}.$
@@ 3.1. Tuning $x_{0}$
    Estimate $\widehat{x_{0}^{\star}}=(\widehat{m}^{2}+\widehat{\sigma^{2}})/\widehat{m}$ where $\widehat{m}$ and $\widehat{\sigma^{2}}$ are the empirical mean and variance of the pilot sample.
@@
    Derive the upper bound of a lower one-sided bootstrap $(1-\alpha)$-confidence interval for $x_{0}^{\min}$ (with $\alpha$ very small); note that $\widehat{x_{0}^{\min}}=\widehat{x_{0}^{\star}}/2$.
@@ 3.2. Setting $p$ and the CPU budget
The reason why we ‘align’ the choice of $p$ with the size of the pilot run $n_{0}$ is discussed next. (Note that $\mathrm{Geometric}(1/(n_{0}+1))$ has mean $n_{0}$.)
Choosing $p$ amounts to choosing the number of variables $X_{i}$ that must be generated, and ultimately the CPU budget of $\hat{f}^{\mathrm{C}}$ since, as we already mentioned, the generation of the $X_{i}$’s often dominates the CPU cost of $\hat{f}^{\mathrm{C}}$.
For both the cycling and simple estimators, we have seen in the previous section that a large pilot run of size $n_{0}$ is often required to choose $x_{0}$ well. It is thus illusory to try to use $\hat{f}^{\mathrm{C}}$ in a low budget regime. The cycling estimator with $\mathbb{E}[R]$ chosen to be the same order as this pilot run size delivers a lower variance estimate with the same order of cost. This makes cycling preferable to the simple estimator.
@@
Figure 1 plots an upper bound on the work-normalised variance, obtained by replacing $\operatorname{var}[\hat{f}]$ with the bounds given in Proposition 4 and 9 for the cycling estimator (Proposition 7 for the simple estimator), as a function of $\mathbb{E}[R]$. We consider $n_{0}=10$ since this is the value used in the experiments in Section 4. When $\mathbb{E}[R]$ gets too small, both bounds diverge at rate $\mathcal{O}\left(1/(\mathbb{E}[R]-\beta^{2})\right)$; when $\mathbb{E}[R]$ tends to infinity, the bound for the cycling estimator grows slowly, at rate $\mathcal{O}(\log\mathbb{E}[R])$, while that for the simple estimator diverges at rate $\mathcal{O}(\mathbb{E}[R])$.

### v3

@@ 1.1. General setting
These conditions ensure that $\hat{f}$ is an unbiased estimator, see the following lemma and Appendix A for a proof.
@@ 1.2. Motivation and contributions
We are particularly interested in the functions $f(x)=\log x$ and $f(x)=1/x$. The former arises in applications such as maximum likelihood estimation where unbiased estimates of the likelihood are available, for example using importance sampling. Using the sum estimator framework, these can then be converted into unbiased estimates of the log-likelihood and its gradient. The function $f(x)=1/x$ arises in Bayesian and frequentist estimation of models with un-normalised likelihoods, that is, when the likelihood involves an intractable normalising constant.
Various unbiased estimators have already been proposed in the literature for such functions; see Section 1.3 for a review. However, in our experience, their performance tends to be very sensitive to ‘tuning’ parameters such as the expansion point $x_{0}$ and the distribution of $R$. This is problematic for users, in particular in situations where one needs to routinely compute unbiased estimates of $f(m)$ at many different $m$ values. This is the case, for instance, in stochastic gradient methods, where at each iteration one needs to estimate the gradient of a different term; see Section 2.6 for details.
Thus, our motivation is to propose a general ‘turnkey’ approach to unbiased estimation which is both robust and easy to automate. Specifically, our contributions are as follows. We derive (Section 2) some results on the general properties of the randomly truncated Taylor-series based estimator, and use them to inform the choice of the truncation random variable $R$. We also discuss in detail the properties of the standard unbiased estimate (4) of $(m/x_{0}-1)^{k}$ and propose an alternative with potentially much smaller variance. In particular, our new estimate uses the whole collection of samples $X_{1},\ldots,X_{r}$ to construct the unbiased estimate $U_{r,k}$ of $(m/x_{0}-1)^{k}$ for $0<k\leq r$. We show the overall variance of the sum estimator will then decrease at a rate $\mathcal{O}\left((\log\mathbb{E}[R])/\mathbb{E}[R]\right)$. In contrast, the variance of the sum estimate with the standard choice (4) for the coefficient estimate does not diminish to zero as $\mathbb{E}[R]$ increases. This difference can be significant for some applications.
The second part of our study (Section 3) focuses on the specific functions $f(x)=\log x$ and $f(x)=1/x$. For these functions, we discuss how to choose tuning parameters, $x_{0}$ and the distribution of $R$, to ensure that the resulting estimator has finite – and hopefully small – variance.
@@ 1.4. Notations and standing assumptions
The expansion point $x_{0}\in\mathbb{R}\setminus\{0\}$ in (1) is fixed initially, and thus so are the $\gamma_{k}$’s, but we discuss how to choose $x_{0}$, for a particular $m$, later on in Section 3.
@@ 2.1. Variance of the general sum estimator
(5) $\mathbb{E}\left[\hat{f}\middle|R=r\right]=\sum_{k=0}^{\infty}\gamma_{k}\frac{\mathds{1}\{r\geq k\}}{\mathbb{P}(R\geq k)}\mathbb{E}[U_{r,k}|R=r]=\sum_{k=0}^{\infty}\gamma_{k}\frac{\mathds{1}\{r\geq k\}}{\mathbb{P}(R\geq k)}\left(\frac{m}{x_{0}}-1\right)^{k}.$
See Appendix A.2 for the verification.
One may decompose its variance, assuming it is finite, through the law of total variance,
@@
See Appendix A.3 for the verification.
The term $\operatorname{var}\left[\mathbb{E}[\hat{f}|R]\right]$ has a simple interpretation, namely, it is the variance of $\hat{f}$ in the ideal case where the $X_{i}$’s and thus the $U_{r,k}$’s have zero variance, that is, $U_{r,k}=(m/x_{0}-1)^{k}$. Hence, this variance component is only due to the random truncation of the Taylor expansion.
@@
For the function $f:\mathbb{R}\rightarrow\mathbb{R}$, argument $m$ and expansion point $x_{0}$, the series expansion in (1) holds. Furthermore, (i) for some $c>0$, $|\gamma_{k}|\leq c$ for all $k=0,1,\ldots$; (ii) $|\gamma_{k}|\exp(\alpha k)\to\infty$ for all $\alpha>0$; (iii) $\beta_{0}:=|m/x_{0}-1|<1$.
@@
See Appendix B.1 for a proof. Assumption 1 is needed to ensure integrability and thus that $\mathbb{E}[\hat{f}|R]$ is well defined. The condition $\beta_{0}<1$ of Assumption 3 and corresponding restriction on $p$ ensures the variance is finite.
@@
Both our functions of interest fulfil Assumption 3: $\gamma_{k}=(-1)^{k-1}/k$ for $f(x)=\log x$, and $\gamma_{k}=(-1)^{k-1}$ for $f(x)=1/x$. A function that does not fulfil Assumption 3 is $f(x)=\exp(x)$ as $\gamma_{k}$ decreases at a super-geometric rate. This makes it possible to consider for $R$ some distribution with tails that are lighter than geometric, for example, the Poisson distribution, which is why we do not considered it further, but again see Papaspiliopoulos, (2011) for a thorough treatment of this function.
Before moving on, we state the expression for the second term of decomposition (6), which we will study in the following sections. See Appendix A.4 for the derivation.
@@ 2.2. Simple estimator
Note that, given $m$ and $\sigma^{2}$, it is always possible to find $x_{0}$ such that $\beta<1$; in fact the minimiser of $\beta$ with respect to $x_{0}$ verifies this condition.
What is unsatisfactory about this result is that the right hand side does not vanish as $p\rightarrow 0$. This is not because the bound is not tight, as shown by the following example whose derivation is given in Appendix B.3. Our numerical experiments suggest that this is true for other functions of interest.
@@ 2.3. Cycling estimator
(12) $U_{r,k}^{\mathrm{C}}:=\frac{1}{r}\left[\prod_{i=1}^{k}\left(\frac{X_{i}}{x_{0}}-1\right)+\prod_{i=2}^{k+1}\left(\frac{X_{i}}{x_{0}}-1\right)+\ldots+\left(\frac{X_{r}}{x_{0}}-1\right)\prod_{i=1}^{k-1}\left(\frac{X_{i}}{x_{0}}-1\right)\right].$
@@
In other words, by taking $\mathbb{E}[R]\rightarrow+\infty$, the cycling estimate recovers, up to a log factor, the standard Monte Carlo rate, i.e. the variance is $\mathcal{O}\left((\log\{\mathbb{E}[R]\})/\mathbb{E}[R]\right)$. Thus, both terms of the variance in (6) now decrease as the mean of $R$ increases.
A direct corollary, through Chebyshev’s inequality, of the above result is that the cycling estimator converges in probability.
@@ 2.4. Minimum variance unbiased estimator
which has minimimum variance among unbiased estimators of its expectation, see Hoeffding, (1948). In spite of the combinatorial number of terms, it is still possible to compute the estimator in $\mathcal{O}(r^{2})$ operations, using a recursion, as pointed out by Sam Power in a personal communication, see Appendix D. However, repeating our variance analysis for the MVUE estimator appears challenging, indeed due to the combinatorial number of terms in (13), and is left for further study. For the moment, we present a numerical comparison in Appendix D, which happens to show cycling and MVUE performing similarly.
@@ 2.5. Relative merits of the simple and cycling estimators
The cycling estimator makes it possible to estimate $f(m)$ as accurately as required, and enjoys, up to a log factor, the same error rate as Monte Carlo estimates. On the other hand, it is more expensive to compute than the simple estimator. Although both estimators require generating $R$ random variables, computing the $U_{R,k}$’s ($0\leq k\leq R$) for the cycling estimator requires $\mathcal{O}(R^{2})$ operations.
@@ 2.6. Unbiased estimators of gradients
and the corresponding ‘sum’ estimator, that is similar to (2) in construction, is
@@
where $R$ is a random variable taking values in $\{1,\ldots\}$; $W_{R,k}=\sum_{r=k}^{\infty}W_{r,k}\mathds{1}\{R=r\}$; $W_{r,k}$ is an unbiased estimate of $(m(\theta)/x_{0}-1)^{k-1}\nabla_{\theta}m(\theta)$ for $r\geq k$; and each and every $W_{r,k}$ is independent of $R$.
As was stated for $\hat{f}$ in Lemma 2, the gradient estimate can be shown to be integrable and unbiased when the estimates $W_{r,k}$ satisfy similar conditions of Assumption 1, that is, with univariates $U_{r,k}$ therein replaced with $W_{r,k}$, and $\gamma_{k}$ replaced with $k\gamma_{k}$. See Appendix A.1 for the proof of unbiasedness of $\hat{f}$ which can be easily adapted to the present gradient case.
We can define a simple estimator for (14), denoted by $\widehat{\nabla_{\theta}f}^{\mathrm{S}}$, by replacing $W_{r,k}$ in (15) with
@@
$W_{r,k}^{\mathrm{C}}:=\frac{1}{r}\Bigg{[}G_{k}\prod_{i=1}^{k-1}\left(\frac{X_{i}}{x_{0}}-1\right)+G_{k+1}\prod_{i=2}^{k}\left(\frac{X_{i}}{x_{0}}-1\right)+\ldots+G_{k-1}\left(\frac{X_{r}}{x_{0}}-1\right)\prod_{i=1}^{k-2}\left(\frac{X_{i}}{x_{0}}-1\right)\Bigg{]}.$
@@ 3.1. Tuning $x_{0}$
    Estimate $x_{0}^{\star}=(m^{2}+\sigma^{2})/m$ as $\widehat{x_{0}^{\star}}=(\widehat{m}^{2}+\widehat{\sigma^{2}})/\widehat{m}$ where $\widehat{m}$ and $\widehat{\sigma^{2}}$ are the empirical mean and variance of the pilot sample.
@@
    Find the upper limit of a lower one-sided bootstrap $(1-\alpha)$-confidence interval for $x_{0}^{\min}$, with $\alpha$ very small; note that $x_{0}^{\min}=x_{0}^{\star}/2$.
@@ 3.2. Setting $p$ and the CPU budget
The reason why we ‘align’ the choice of $p$ with the size of the pilot run $n_{0}$ is discussed next. Note that $\mathrm{Geometric}(1/(n_{0}+1))$ has mean $n_{0}$. The resulting procedure is summarised in Algorithm 1 for the cycling estimator. The simple estimator is obtained by replacing $U_{r,k}^{\mathrm{C}}$ with $U_{r,k}^{\mathrm{S}}$.
1:  Inputs: number of pilot runs $n_{0}$, a mechanism to sample $X_{i}$ with mean $m$, confidence level $\alpha$, $(\gamma_{k})_{k\geq 0}$.
2:  Tuning:
3:  Sample $X_{1},\dots X_{n_{0}}$ and compute empirical mean $\widehat{m}$ and variance $\widehat{\sigma^{2}}$.
4:  Estimate $\widehat{x_{0}^{\star}}=(\widehat{m}^{2}+\widehat{\sigma^{2}})/\widehat{m}$.
5:  Build a one-sided bootstrap $(1-\alpha)$-confidence interval $(-\infty,u]$, for $\sigma^{2}/(2m)+m/2$.
6:  Set $x_{0}=\max\{u,\widehat{x_{0}^{\star}}\}$ and $\widehat{\beta^{2}}=\widehat{\sigma}^{2}/x_{0}^{2}+(\widehat{m}/x_{0}-1)^{2}$.
7:  Set $p=\min\left(1-\widehat{\beta^{2}},\frac{1}{n_{0}+1}\right)$.
8:  Estimator:
9:  Sample $R\sim\mathrm{Geometric}(p)$ and, given $R=r$, sample $X_{1},\dots,X_{r}$.
10:  Compute $U_{r,k}^{\mathrm{C}}$ as in (12).
11:  Output: $\hat{f}^{\mathrm{C}}=\sum_{k=0}^{r}\frac{\gamma_{k}U_{r,k}^{\mathrm{C}}}{(1-p)^{k}}$.
Algorithm 1 Cycling Taylor-based sum estimator
Choosing $p$ amounts to choosing the number of variables $X_{i}$ that must be generated, and ultimately the CPU budget of $\hat{f}^{\mathrm{C}}$ since, as we already mentioned, the generation of the $X_{i}$’s often dominates the CPU cost of $\hat{f}^{\mathrm{C}}$. For both the cycling and simple estimators, we have seen in the previous section that a large pilot run of size $n_{0}$ is often required to choose $x_{0}$ well. It is thus illusory to try to use $\hat{f}^{\mathrm{C}}$ in a low budget regime.
@@
Figure 1 plots an upper bound on the work-normalised variance, obtained by replacing $\operatorname{var}[\hat{f}]$ with the bounds given in Proposition 4 and 9 for the cycling estimator (Proposition 7 for the simple estimator) in the case $x_{0}=x_{0}^{\star}$, as a function of $\mathbb{E}[R]$. We consider $n_{0}=10$ since this is the value used in the experiments in Section 4. When $\mathbb{E}[R]$ gets too small, both bounds diverge at rate $\mathcal{O}\left(1/(\mathbb{E}[R]-\beta^{2})\right)$; when $\mathbb{E}[R]$ tends to infinity, the bound for the cycling estimator grows slowly, at rate $\mathcal{O}(\log\mathbb{E}[R])$, while that for the simple estimator diverges at rate $\mathcal{O}(\mathbb{E}[R])$.
[…]